OpenStudy (anonymous):

Verify if \[x+-\frac{3}{2}(x+y-2)+\ln|x+y-2|=c\] is a solution to \[(x+y) dx + (3x+3y-4)dy = 0\]

OpenStudy (anonymous):

Or you could start with the latter equation and see if it ends up with the former equation if it's too hard. I substituted y with \[t=a_1x+b_1y => y = \frac{t-a_1x}{b_1}\] and it follows that\[dy=\frac{dt-a_1dx}{b_1}\]

OpenStudy (anonymous):

diff equation can be done by sub \(z=x+y\)

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{x+y}{4-3(x+y)}\]

OpenStudy (anonymous):

Well I have to solve it in a specific way. It's a homogeneous differential equation where M(x,y) and N(x,y) are linear but not homogeneous. The equations in M(x,y) and N(x,y) actually represent equations of lines, which in this case are parallel.

OpenStudy (anonymous):

sorry i was out so u want to make it homogeneous with mapping \((x,y)\) to \((X,Y)\) so that the diff equation will be homogeneous in tems of \(X\) and \(Y\)

OpenStudy (anonymous):

u can show that the sub \(X=x-x_0\) and \(Y=y-y_0\) will change the nonhomogeneous diff equation\[\frac{dy}{dx}=\frac{ax+by+c}{ex+fy+g}\]to homogeneous diff equation\[\frac{dY}{dX}=\frac{aX+bY}{eX+fY}\]where \((x_0,y_0)\) is solution of system\[ax+by+c=0\]\[ex+fy+g=0\]